What similar pieces of information do you think you would get from a dot blot analysis compared to a western blot analysis?

Think about the two methods (the western and dot blot) how they differ, what unique benefits might each technique present? When might one be more useful than the other?

You have a stock of EDTA (0.1M) and you need 100ml of 3mM EDTA. Explain how you would make up this solution in the lab.

A volume of serum sample is diluted with equal amounts of buffer (1:1). Following this, a series of five dilutions is made of this first dilution by diluting it 1/10, rediluting 1/10, and then three times more, each resulting solution then being a 1/10 dilution of the previous one in the series. What is the concentration of serum in each solution made?
Answer #1

1. Dot not and western blot both are used to study protein of interest or protein studies

Both utilise antibody mediated detection

2 . 0.1 M or 100 mM

Desired concentration : 3 mM

desired volume : 100 ml

V1N1 = V2N2

V1*100 = 100*3

V1= 300/100

V1 = 3 ml

Take 3 ml from 0.1 M stock and add 97 ml of distilled water

3. Serum concentration after diluting in 1:1 buffer is 50

Upon diluting 1/10 = 5

2nd 1/10 dilution = 0.5

3rd 1/10 dilution = 0.05

4th 1/10 dilution = 0.005

5th 1/10 dilution = 0.0005 or 5 *10-4